OpenStudy (anonymous):

Simplify: ( 49/36 )^-1/2

OpenStudy (nenadmatematika):

\[=1/\sqrt{49/36}\]=1/(7/6)=6/7

OpenStudy (precal):

the negative power changes it to (36/49)^(1/2) I am by the way assuming that it is to the (-1/2) power \[\sqrt{36}/\sqrt{49}\] 6/7